OpenStudy (anonymous):

What is the equation of the line with a slope of 2 and a y-intercept of −4? y = mx + b m = 2 and b = −4 y = 2x − 4 I understand the concept but how do I know which kind of equation it asks for, standard form or slope intercept? For this question I know it's slope intercept but I ask for future reference. @Sheraz12345 @ganeshie8

OpenStudy (anonymous):

You are absolutely right ;o

OpenStudy (anonymous):

well when you have two points, use slope intercept. when it says AX+BY=C use standard

OpenStudy (anonymous):

I know, but what I'm asking is how do I know that the question is asking for either one?

OpenStudy (sheraz12345):

@LydiannAquino what exactly is standard form.... can u explain meh?

OpenStudy (anonymous):

y=mx+b

OpenStudy (sheraz12345):

and slope intercept form ?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (haseeb96):

actually u have slope which is represented by m =2 and the y intercept which is represented by ''b'' so the slope intercept form formula is y=mx+b so just put and get the answer which is y= 2x-4 but here it is asking about the equation line so if the question said to u write the equation of line in slope intercept form then u will write but if not the convert it into standard form that is 2x -y-4=0 is the equation of line

OpenStudy (anonymous):

Ok, Thank You.

OpenStudy (haseeb96):

have u got it ?

OpenStudy (anonymous):

Yes.